### v2

# Regular graphs are universally 3-edge-weightable

Abstract.
A graph is universally $k$-edge-weightable if for every $k$-element set $Q\subset\mathbb{R}$, it admits a proper $Q$-edge weighting. The settled 1-2-3 conjecture implies that for any arithmetic progression $\{a,b,c\}$, every nice regular graph has a proper $\{a,b,c\}$-edge weighting. We prove that this remains valid for all 3-element set $\{a,b,c\}$ with $c-b\neq b-a$. Consequently, every nice regular graph is universally $3$-edge-weightable.
Key words and phrases:
edge weighting, regular graph, 1-2-3 conjecture, universal 3-edge-weightability
2020 Mathematics Subject Classification:
05C15,05C78
Kecai Deng is the corresponding author.
  Kecai Deng is partially supported by NSFC (No. 11701195), and Fundamental Research Funds for the Central Universities (No. ZQN-904).
1. Introduction
All graphs considered are simple, finite, and undirected. For notation or terminology not defined here, we refer to [1]. A graph is nice if it contains no $K_{2}$ component. For $G=(V,E)$ and $Q\subset\mathbb{R}$, a $Q$-edge weighting is a mapping $w:E\to Q$. When $Q=\{1,2,\ldots,k\}$, it is called a $k$-edge weighting. The weighted degree of $v$ is $d_{w}(v)=\sum_{e\ni v}w(e)$. An edge $uv$ is a conflict if $d_{w}(u)=d_{w}(v)$; $w$ is proper if it has no conflicts.
The 1-2-3 conjecture, posed by Karoński, Łuczak, and Thomason [2], asserts that every nice graph has a proper $\{1,2,3\}$-edge weighting. This conjecture is a local version of the classical–and still active–irregularity strength problem [3, 4, 6, 7, 5, 8, 9]. Research on the 1-2-3 conjecture has progressed along several fronts. Upper bounds for general graphs were successively lowered: from 30 [10], to 16 [11], 13 [12], 5 [13], and finally to 3 [14] after an intermediate improvement to 4 [15]. For specific classes, it was verified for dense graphs [16], $d$-regular graphs with large $d$ [17], graphs with $\delta=\Omega(\log\Delta)$ [18], and shown to hold asymptotically almost surely for random graphs [11]. Complexity results include NP-completeness for deciding proper $\{1,2\}$-edge weightings [19] and a polynomial-time algorithm for bipartite graphs [20]. The conjecture was affirmatively resolved by Keusch [14], leading to the 1-2-3 theorem. For more 1-2-3 conjecture related problems, refer to [21, 22, 23, 24, 25, 26, 27] and the references cited there.
Various strengthenings of the 1-2-3 conjecture have been studied. The list version [28] asks whether every nice graph is edge-weight $3$-choosable, i.e. for any assignment of $3$-element lists $L(e)$ to edges, there exists a proper edge weighting $w$ with $w(e)\in L(e)$ for all $e$. Wong and Zhu [29] conjectured a stronger property called $(1,3)$-choosability, where each vertex receives a fixed weight (from a $1$-element list) and each edge receives a weight from a $3$-element list; the weighted degree of a vertex is then $d_{w}(v)=w(v)+\sum_{e\ni v}w(e)$. These properties form a clear hierarchy:
$(1,3)\text{-choosable}\Rightarrow\text{edge-weight 3-choosable}\Rightarrow\text{the 1-2-3 theorem}.$
Progress on the strongest property, $(1,3)$-choosability, has been steady: Cao [30] proved that every nice graph is $(1,17)$-choosable, improved by Zhu [31] to $(1,5)$-choosability. Consequently, every nice graph is edge-weight $5$-choosable. However, it remains open whether every nice graph is $(1,3)$-choosable, or even edge-weight $3$-choosable. Partial results show that several graph classes are $(1,3)$-choosable, including complete graphs and complete bipartite graphs [28], generalized Halin graphs [32], graphs with maximum average degree less than $\frac{11}{4}$ [33], dense Eulerian graphs [34], and Cartesian products of paths and cycles [35].
The 1-2-3 theorem guarantees a proper $\{1,2,3\}$-edge weighting for every nice graph. A natural weakening of the unsolved $(1,3)$-choosability conjecture and edge-weight $3$-choosability conjecture, is to ask whether this extends to any three-element set $Q\subset\mathbb{R}$, not just $\{1,2,3\}$. We call a graph universally $3$-edge-weightable if it admits a proper $Q$-edge weighting for every such $Q$.
For regular graphs, affine transformation from the 1-2-3 theorem immediately gives a proper $\{a,b,c\}$-edge weighting for any arithmetic progression $\{a,b,c\}$ (where ba=cbb-a=c-b). The non-arithmetic case, however, requires genuinely new constructions. This paper provides the first such construction, which implies that every regular graph is universally $3$-edge-weightable.
Theorem 1.
For any positive real numbers $d_{1},d_{2}$ with $0<d_{1}<d_{2}$, every $k$-regular graph with $k\geq 3$ admits a proper $\{-d_{1},0,d_{2}\}$.
Corollary 2.
Every $k$-regular graph $(k\geq 2)$ is universally $3$-edge-weightable.
Proof.
For $k=2$, a $2$-regular graph consists of disjoint cycles. Graphs with maximum average degree less than $\frac{11}{4}$ are known to be $(1,3)$-choosable [33], implying the existence of a proper $\{a,b,c\}$-edge weighting. Thus we may assume $k\geq 3$.
It suffices to show that $G$ admits a proper $\{-(b-a),0,c-b\}$-edge weighting $w^{\prime}$. Indeed, if we define $w=w^{\prime}+b$ (that is, we add $b$ to every edge weight), the regularity of $G$ ensures that $w$ is a proper $\{a,b,c\}$-edge weighting.
Consider three cases. If ba<cbb-a<c-b, Theorem 1 directly provides a proper $\{-(b-a),0,c-b\}$-edge weighting. If ba=cbb-a=c-b, Keusch’s theorem [14] gives a proper $\{1,2,3\}$-edge weighting $w^{\prime\prime}$ of $G$. Setting $w^{\prime}=(b-a)(w^{\prime\prime}-2)$ yields a proper $\{-(b-a),0,c-b\}$-edge weighting. Finally, if ba>cbb-a>c-b, Theorem 1 supplies a proper $\{-(c-b),0,b-a\}$-edge weighting $w^{\prime\prime\prime}$; then $w^{\prime}=-w^{\prime\prime\prime}$ is the desired weighting. In all cases the required weighting exists, proving the corollary. ∎
This settles the regular case completely. Whether universal $3$-edge-weightability holds for all nice graphs remains open:
Problem 1.1 (Universal 3-edge-weightability).
Is every nice graph universally $3$-edge-weightable?
Theorem 1 is proved in Section 2.
2. Proof of Theorem 1
We will use the following three lemmas.
Lemma 1.
[1] Each $i$-degenerate graph $(i\geq 1)$ is $(i+1)$-colorable.
Lemma 2.
[36] Let $G=(V,E)$ be a graph, $I_{0}$ be a maximum independent set of $G$, and $I_{1}$ be an independent set of $G[V\setminus I_{0}]$. Then the bipartite graph $G[I_{0},I_{1}]$ admits a matching saturating $I_{1}$.
Lemma 3.
Let $G=(I_{0}\cup I_{1},E)$ be a bipartite graph with $d(v)\geq 3$ for each $v\in I_{1}$; let $d_{1}$ and $d_{2}$ be two real numbers with $0<d_{1}<d_{2}$. Then $G$ admits a proper $\{-d_{1},0,d_{2}\}$-edge weighting.
Proof.
We construct an edge weighting $w$ such that every vertex in $I_{1}$ attains a weighted degree in $\{-3d_{1},d_{2}\}$ and each one in $I_{0}$ has weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
Initially, we assign weight 0 to every edge. Let $I_{1}=\{x_{i}:i=1,2,\ldots,p\}$.
First, label three of the edges incident to $x_{1}$ with $d_{1}$. Then $d_{w}(x_{1})=-3d_{1}$, and for each $y\in I_{0}$, $d_{w}(y)\in\{0,-d_{1}\}\subseteq(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
Second, for $i=2,3,\ldots,p$, suppose:
(i) for each $j\geq i$, no edge incident to $x_{j}$ is weighted;
(ii) for each vertex $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$;
(iii) for each $j<i$, $d_{w}(x_{j})\in\{-3d_{1},d_{2}\}$.
Case 1: If $x_{i}$ has a neighbor $y^{\ast}$ with $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$, let $w(x_{i}y^{\ast})=d_{2}$. After this assignment, $d_{w}(x_{i})=d_{2}$, $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, and each $y\in I_{0}\setminus\{y^{\ast}\}$ still satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Thus, conditions (i)-(iii) are maintained for $i+1$.
Case 2: If each neighbor of $x_{i}$ has a weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$, let $y_{1},y_{2}$ and $y_{3}$ be three of them. Set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. Then $d_{w}(x_{i})=-3d_{1}$. For each $j=1,2,3$, $d_{w}(y_{j})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. What is more, each $y\in I_{0}\setminus\{y_{1},y_{2},y_{3}\}$ still satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Thus, conditions (i)-(iii) hold for $i+1$.
Consequently, each vertex in $I_{0}$ attains a weighted degree in $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$, while each one in $I_{1}$ has weighted degree in $\{-3d_{1},d_{2}\}$, and we obtain a proper $\{-d_{1},0,d_{2}\}$-edge weighting. ∎
Proof of Theorem 1.
Let $G=(V,E)$ be a $k$-regular graph with $k\geq 3$.
Partition. We partition $V$ into independent sets as follows. Choose a maximum independent set $I_{0}$ in $G$. For $i\geq 1$, let $I_{i}$ be a maximum independent set in $G_{i}$, where $G_{i}\triangleq G[V\setminus(\cup_{j<i}I_{j})]$. Let $\ell$ be the maximum integer such that $I_{\ell}\neq\emptyset$. Then $I_{\ell}$ is an independent set, and $1\leq\ell\leq k+1$ since $G$ is nonempty and each vertex in $I_{\ell}$ has a neighbor in $I_{i}$ for each $i<\ell$. If $\ell=1$, the conclusion holds from Lemma 3. So suppose $2\leq\ell\leq k+1$.
We will construct a proper $\{-d_{1},0,d_{2}\}$-edge weighting $w$ in two phases. In the weighting process, for $i=1,2,\ldots,\ell,$ a vertex $v$ in $I_{i}$ is called:
  • Type I if $d_{w}(v)=(i-1)d_{2}$, and $v$ is incident to exactly i1i-1 edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$;
  • Type II if $d_{w}(v)=id_{2}-d_{1}$, and $v$ is incident to exactly $i$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$ and one edge of weight $d_{1}$ connecting $v$ to $I_{0}$;
  • Type III if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=id_{2}-2d_{1}$, and $v$ is incident to exactly $i$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$ and two edges of weight $d_{1}$: one connecting $v$ to $I_{0}$ and the other connecting $v$ to a vertex in $I_{1}$;
  • hungry if $v\in I_{i}$ for some $i\geq 2$, $d_{w}(v)=pd_{2}$ for some $p\leq i-2$ and $v$ is incident to exactly $p$ edges of weight $d_{2}$ connecting $v$ to $\cup_{j\geq 1}I_{j}$.
Goal. We will make sure $w$ finally satisfies the following conditions:
  • (1)
    for each vertex $v$ in $I_{i}$ with $i\geq 2$, $d_{w}(v)\in\{(i-1)d_{2},id_{2}-2d_{1},id_{2}-d_{1}\}$ (that is, $v$ is Type I, II or III);
  • (2)
    for each $x$ in $I_{1}$, $d_{w}(x)\in\{-3d_{1},-d_{1},0,d_{2}-d_{1}\}$;
  • (3)
    for each $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})$;
  • (4)
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-3d_{1}$, $x$ has at least two neighbors in $I_{0}$, and every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$;
  • (5)
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=-d_{1}$, $x$ has a unique neighbor $y_{x}$ in $I_{0}$, and $d_{w}(y_{x})=0$;
  • (6)
    for each vertex $x$ in $I_{1}$ with $d_{w}(x)=0$, every neighbor $y$ of $x$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{0\}$.
Note that in (1), $\max\{(i-1)d_{2},id_{2}-2d_{1}\}<id_{2}-d_{1}$ because $d_{1}<d_{2}$. Moreover, for $1\leq i<j$, one has $id_{2}-d_{1}<\min\{(j-1)d_{2},jd_{2}-2d_{1}\}$. So (1)-(3) ensure there is no conflict in $G[\cup_{i\geq 1}I_{i}]$, and there is no conflict between $I_{0}$ and $\cup_{i\geq 2}I_{i}$. Clearly, (3)-(6) ensure there is no conflict between $I_{0}$ and $I_{1}$.
The construction proceeds in two phases. Phase 1 ensures condition (1)-(3) hold; Phase 2 ensures conditions (4)-(6) hold, with (1)-(3) remaining valid.
In the beginning, every edge is weighted with 0.
Phase 1: Initial weighting
In this phase, we will make sure:
  • (1’)
    every vertex in $\cup_{i\geq 1}I_{i}$ is Type I or II;
  • (2’)
    for each $y$ in $I_{0}$, $d_{w}(y)\in(-\infty,0]$;
  • (3’)
    no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$.
Clearly, the validity of conditions (1’)-(3’) guarantees that (1)-(3) hold.
Step 1 By Lemma 2, there exists a matching $M_{\ell-1,\ell}$ in the bipartite graph $G[I_{\ell-1},I_{\ell}]$ that saturates $I_{l}$. Weight each edge in $M_{\ell-1,\ell}$ with $d_{2}$. Then when $\ell=2$, each vertex in $I_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$); when $\ell\geq 3$, each vertex in $I_{\ell}$ is hungry (has weighted degree $d_{2}\leq(\ell-2)d_{2}$). Let $I^{\ast}_{\ell-1}$ be the set of vertices in $I_{\ell-1}$ that are saturated by $M_{\ell-1,\ell}$.
  • When $\ell=2$, by Lemma 2, there exists a matching $M^{\ast}_{0,1}$ in the bipartite graph $G[I_{0},I^{\ast}_{1}]$ that saturates $I^{\ast}_{1}$; weight each edge in $M^{\ast}_{0,1}$ with $d_{1}$, which makes each vertex in $I^{\ast}_{1}$ Type II (has weighted degree $d_{2}-d_{1}$), and each one in $I_{1}\setminus I^{\ast}_{1}$ Type I (has weighted degree $0=(1-1)d_{2}$). Then (1’) and (2’) already hold in this case.
  • When $\ell=3$, each vertex in $I^{\ast}_{2}$ is Type I (has weighted degree $d_{2}=(2-1)d_{2}$), each vertex in $I_{2}\setminus I^{\ast}_{2}$ is hungry (has weighted degree $0\leq(2-2)d_{2}$), and each (hungry) vertex $u$ in $I_{3}$ is incident to an edge of weight $d_{2}$ whose other endpoint lies in $I_{2}$.
  • When $\ell\geq 4$, each vertex in $I_{\ell-1}$ is hungry (has weighted degree at most $d_{2}\leq[(\ell-1)-2]d_{2}$), and each (hungry) vertex $u$ in $I_{\ell}$ is incident to an edge of weight $d_{2}$ whose other endpoint lies in $I_{\ell-1}$.
Step 2 When $\ell\geq 3$, we process the indices $i=\ell-2,\ell-3,\ldots,1$ in decreasing order. Right before proceeding $i$, we assume:
  • (A)
    no edge incident to $\cup_{1\leq j\leq i}I_{j}$ has been weighted;
  • (B)
    for each $\alpha\geq i+1$, every vertex in $I_{\alpha}$ is Type I, Type II or hungry;
  • (C)
    for $\alpha\geq i+2$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta=i+1,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
(The starting point of this inductive hypothesis holds true after Step 1.)
Let $E_{i+1}$ be the set of edges with weight 0 in the subgraph $G[\cup_{j\geq i+1}I_{j}]$. If $E_{i+1}\neq\emptyset$, we process the edges in $E_{i+1}$ one by one. When considering an edge $e$, if both endpoints are hungry, we set $w(e)=d_{2}$; otherwise we leave $e$ unweighted. After processing all edges in $E_{i+1}$, let $U_{i+1}$ be the set of hungry vertices in $\cup_{j\geq i+1}I_{j}$. By construction, if an edge has both endpoints in $U_{i+1}$, then it has already been weighted with $d_{2}$; hence the induced subgraph $G[U_{i+1}]$ (if nonempty) has all edges weighted with $d_{2}$.
Take an arbitrary vertex $u\in U_{i+1}$, and suppose $u\in I_{\alpha}$ for some $\alpha=i+1,\ldots,\ell$. Since $u$ is hungry, $d_{w}(u)=jd_{2}$ for some $j\leq\alpha-2$. By assumption (C), for each $\beta=i+1,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$. There are $(\alpha-1)-(i+1)+1=\alpha-i-1$ such indices $\beta$. Thus at most $(\alpha-2)-(\alpha-i-1)=i-1$ edges of weight $d_{2}$ incident to $u$ can have their other endpoints in $\cup_{j\geq\alpha+1}I_{j}$. Recall that $G[U_{i+1}]$ has all edges weighted with $d_{2}$. So within the subgraph $G[U_{i+1}]$, the vertex $u$ has at most i1i-1 neighbors in $U_{i+1}\cap(\cup_{j\geq\alpha+1}I_{j})$. Since $I_{\alpha}$ is an independent set and $u$ is arbitrary, it follows that $G[U_{i+1}]$ is (i1)(i-1)-degenerate. By Lemma 1, $U_{i+1}$ can be partitioned into at most $i$ independent sets in $G[U_{i+1}]$. Each such independent set is also independent in $G$. Hence we can write $U_{i+1}=J_{1}\oplus\cdots\oplus J_{r}$ with $r\leq i$, where each $J_{t}$ is independent in $G$.
By Lemma 2, for each $t=1,2,\ldots,r$, there exists a matching $M^{\prime}_{i,t}$ in the bipartite graph $G[I_{i},J_{t}]$ that saturates $J_{t}$. For every edge $e$ in $\bigcup_{t=1}^{r}M^{\prime}_{i,t}$, set $w(e)=d_{2}$. At this point, each vertex $v\in I_{i}$ is incident to at most $r$ edges of weight $d_{2}$ from these matchings, so $d_{w}(v)=jd_{2}$ for some $j\leq r\leq i$.
Let $\overline{I}_{i}\subseteq I_{i}$ be the set of vertices with weighted degree exactly $id_{2}$. Applying Lemma 2 again, there exists a matching $\overline{M}_{0,i}$ in $G[I_{0},\overline{I}_{i}]$ that saturates $\overline{I}_{i}$. For each edge $e\in\overline{M}_{0,i}$, set $w(e)=-d_{1}$. Then each vertex in $\overline{I}_{i}$ becomes Type II, and each one in $I_{i}\setminus\overline{I}_{i}$ is either Type I or hungry.
After these assignments, the following hold:
  • (A’)
    no edge incident to $\cup_{1\leq j\leq i-1}I_{j}$ has been weighted;
  • (B’)
    for each $\alpha\geq i$, every vertex in $I_{\alpha}$ is Type I, Type II or hungry;
  • (C’)
    for $\alpha\geq i+1$, if $I_{\alpha}$ has hungry vertices, then: for every hungry vertex $u$ in $I_{\alpha}$ and every $\beta=i,\ldots,\alpha-1$, $u$ is incident to at least one edge of weight $d_{2}$ whose other endpoint lies in $I_{\beta}$.
Indeed, (B’) holds because: if $v\in\cup_{j\geq i+1}I_{j}$ was already Type I or II after step $i+1$, it remains Type I or II; if $u\in U_{i+1}$ received an edge of weight $d_{2}$ from some $M^{\prime}_{i,t}$, it now becomes Type I or remains hungry (the extra $d_{1}$ edges to $I_{0}$ do not affect the condition for vertices in $U_{i+1}$); if $v\in I_{i}$, then either it received at most i1i-1 edges of weight $d_{2}$ from matchings $M^{\prime}_{i,t}$’s (in which case it becomes Type I or remains hungry), or received exactly $i$ edges of weight $d_{2}$ and exactly one edge of weight $d_{1}$ from $\overline{M}_{0,i}$ (in which case it becomes Type II).
Proceeding in this way for $i=\ell-2,\ell-3,\ldots,1$, we obtain at the end that every vertex in $\cup_{\alpha\geq 1}I_{\alpha}$ is Type I or II. Since at the end, on one hand, each $v$ in $I_{1}$ has weighted degree $d_{2}-d_{1}$ (Type II) or 0 (Type I). On the other hand, if some $u$ in $I_{\alpha}$ ($\alpha\geq 2$) is hungry, it should satisfies (C’): $u$ is incident to at least one edge of weight $d_{2}$ with other endpoint in $I_{\beta}$ for each $\beta=1,\ldots,\alpha-1$ (at the end $i=1$). That is, $d_{w}(u)\geq(\alpha-1)d_{2}$ which implies that $u$ is not hungry, a contradiction. Consequently, condition (1’) holds. Moreover, every $y$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,0]$ at this stage and (2’) holds.
For (3’), if there exists a Type I vertex $v$ in $I_{1}$ that has a Type I neighbor $u$ in some $I_{i}$ ($i\geq 2$), let $e_{v}$ and $e_{u}$ be the edges joining $v$ and $u$ to $I_{0}$, respectively (recall that each vertex in $\cup_{j\geq 1}I_{j}$ has a neighbor in $I_{0}$ since $I_{0}$ is a maximum independent set). Clearly, $e_{v},e_{u}$ and $uv$ have not been weighted since $v$ and $u$ are both Type I (moreover, $v$ is in $I_{1}$). Set $w(e_{v})=w(e_{u})=-d_{1}$ and set $w(uv)=d_{2}$. Then both $u$ and $v$ become Type II (so condition (1’) remains valid), and every vertex in $I_{0}$ still has weighted degree at most 0 (so condition (2’) remains valid). Repeat this step until no Type I vertex in $I_{1}$ has a Type I neighbor in $\cup_{i\geq 2}I_{i}$, and (3’) holds finally.
Phase 2: Conflict resolution
For $i=0,1$, let $I_{i,0}=\{v\in I_{i}:d_{w}(v)=0\}.$ By our construction, all edges incident to vertices in $I_{0,0}\cup I_{1,0}$ currently have weight 0. After Phase 1, by (1’) and (2’), the only possible conflicts in $G$ are those belonging to the bipartite subgraph $G[I_{0,0},I_{1,0}]$. If $G[I_{0,0},I_{1,0}]$ is empty, we are done. Assume therefore that it is nonempty.
For each $x\in I_{1,0}$ with $N(x)\cap I_{0,0}\neq\emptyset$, choose an arbitrary vertex from $N(x)\cap I_{0}$ and denote it by $y_{x}$. For $i\geq 1$, let $X_{i}=\{x\in I_{1,0}:|N(x)\cap I_{0}|=i\}$ and $X_{i^{+}}=\cup_{j\geq i}X_{j}$.
Now consider the bipartite subgraph $H^{\prime}:=G[X_{1}\cup X_{2},\cup_{i\geq 2}I_{i}]$. On one hand, for any $x\in X_{1}\cup X_{2}$, the vertex $x$ has at least k2k-2 Type II neighbors in $\bigcup_{i\geq 2}I_{i}$. This holds because, by conditions (1’) and (3’), every neighbor of $x$ either lies in $I_{0}$ or is a Type II vertex in $\bigcup_{i\geq 2}I_{i}$. On the other hand, consider a Type II vertex $z\in I_{i}$ with $i\geq 2$. By definition, $z$ is incident to exactly $i$ edges of weight $d_{2}$ and one edge of weight $d_{1}$. Since all edges incident to vertices in $X_{1}\cup X_{2}$ currently carry weight 0, any edge of weight $d_{2}$ or $d_{1}$ incident to $z$ must have its other endpoint outside $X_{1}\cup X_{2}$. Thus $z$ can be adjacent to at most k2k-2 vertices in $X_{1}\cup X_{2}$. Thus, in $H^{\prime}$, every vertex in $X_{1}\cup X_{2}$ has degree at least k2k-2, and every vertex in $\cup_{i\geq 2}I_{i}$ has degree at most k3k-3. By Hall’s theorem, there exists a matching $M$ in $H^{\prime}$ that saturates $X_{1}\cup X_{2}$. Write $M=M_{1}\oplus M_{2}$, where $M_{i}$ saturates $X_{i}$ for $i=1,2$. For each $x\in X_{1}\cup X_{2}$, let $z_{x}$ be the (Type II) vertex in $\cup_{i\geq 2}I_{i}$ such that $xz_{x}\in M$.
We will resolve the conflicts involving $X_{2^{+}}$ and $X_{1}$ separately.
1. Resolving conflicts incident to $X_{2^{+}}$.
Let $X^{\ast}_{2^{+}}\subseteq X_{2^{+}}$ be the set of vertices that have a neighbor in $I_{0,0}$. If $X^{\ast}_{2^{+}}=\emptyset$, then we are done. Suppose $X^{\ast}_{2^{+}}\neq\emptyset$.
We select a vertex $x_{i}$ from $X^{\ast}_{2^{+}}$ (arbitrarily for $i=1$, and from the updated $X^{\ast}_{2^{+}}$ for $i>1$) and handle it according to the cases below. After processing $x_{i}$, it will be removed from $I_{1,0}$ (since its weighted degree will become non-zero), and consequently from $X_{2^{+}}$ and $X^{\ast}_{2^{+}}$ (this process will gradually eliminates vertices from $X_{2^{+}}$ until $X^{\ast}_{2^{+}}$ becomes empty). We then update $I_{1,0}$, $X_{2^{+}}$, $X^{\ast}_{2^{+}}$ accordingly.
Before processing $x_{i}$, we assume:
  • (i)
    no edge incident to unprocessed vertices in $X_{2^{+}}$ has been weighted;
  • (ii)
    for every $y\in I_{0}$, $d_{w}(y)\in(-\infty,d_{2}-d_{1})$;
  • (iii)
    each previously processed vertex $x_{j}$ ($j<i$) satisfies one of the following conditions:
    • $d_{w}(x_{j})=d_{2}-d_{1}$;
    • $d_{w}(x_{j})=-3d_{1}$ and every neighbor $y$ of $x_{j}$ in $I_{0}$ satisfies $d_{w}(y)\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$.
(For $i=1$, conditions (i) and (ii) hold; no vertex has been processed and so condition (iii) holds vacuously.)
We distinguish two cases.
Case 1: If there exists $y^{\ast}\in N(x_{i})\cap I_{0}$ such that $d_{w}(y^{\ast})\in(-\infty,-d_{1})\setminus\{-d_{2}-3d_{1}\}$. Clearly, $y^{\ast}\neq y_{x_{i}}$, since by the definition of $y_{x_{i}}$, we have $d_{w}(y_{x_{i}})=0$ at this stage. Set $w(x_{i}y^{\ast})=d_{2}$ and $w(x_{i}y_{x_{i}})=-d_{1}$. After these assignments, $d_{w}(x_{i})=d_{2}-d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$ and $d_{w}(y^{\ast})\in(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{x_{i}},y^{\ast}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged. Thus, conditions (i)-(iii) continue to hold for $i+1$.
Case 2: If every vertex in $N(x_{i})\cap I_{0}$ has weighted degree in $d_{1},d_{2}-d_{1})\cup\{-d_{2}-3d_{1}\}$.
Subcase 2.1: If $x_{i}\in X_{3^{+}}$. Let $y_{1},y_{2},y_{3}\in N(x_{i})\cap I_{0}$ and set $w(x_{i}y_{1})=w(x_{i}y_{2})=w(x_{i}y_{3})=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{1}),d_{w}(y_{2}),d_{w}(y_{3})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$. Then every vertex in $N(x_{i})\cap I_{0}$ has weighted degree in $2d_{1},d_{2}-d_{1})\cup\{-d_{2}-4d_{1},-d_{2}-3d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{1},y_{2},y_{3}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged. Thus, conditions (i)-(iii) continue to hold for $i+1$.
Subcase 2.2: If $x_{i}\in X_{2}$. Assume $N(x_{i})\cap I_{0}=\{y_{x_{i}},y^{\prime}\}$ and $z_{x_{i}}\in I_{q}$ for some $q\geq 2$. Recall that $z_{x_{i}}$ is Type II and so $d_{w}(z_{x_{i}})=qd_{2}-d_{1}$. Now set $w(x_{i}y_{x_{i}})=w(x_{i}y^{\prime})=w(x_{i}z_{x_{i}})=-d_{1}$. This gives $d_{w}(x_{i})=-3d_{1}$, while $d_{w}(y_{x_{i}})=-d_{1}$, $d_{w}(y^{\prime})\in[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$ and $d_{w}(z_{x_{i}})=qd_{2}-2d_{1}$ (that is, $z_{x_{i}}$ becomes Type III and so condition (1) remains valid). Then $\{d_{w}(y_{x_{i}}),d_{w}(y^{\prime})\}\subseteq[-2d_{1},d_{2}-2d_{1})\cup\{-d_{2}-4d_{1}\}$, which is a subset of $(-\infty,d_{2}-d_{1})\setminus\{-3d_{1}\}$. Moreover, the weighted degree of each vertex in $(I_{0}\setminus\{y_{x_{i}},y^{\prime}\})\cup(X_{2^{+}}\setminus\{x_{i}\})$ remains unchanged. Thus, conditions (i)-(iii) continue to hold for $i+1$.
Remark on the dynamics of $I_{0,0}$: In Case 1, it is possible that $d_{w}(y^{\ast})$ becomes 0 after adding $d_{2}$ (if originally $d_{w}(y^{\ast})=-d_{2}$), placing $y^{\ast}$ back into $I_{0,0}$. However, by assumption (i), all edges incident to $\{x_{j}:j>i\}$ are still unweighted, so any new conflict involving $y^{\ast}$ and $x_{j}$ ($j>i$) will be resolved when $x_{j}$ is processed. Moreover, $y^{\ast}$ cannot conflict with already processed vertices $x_{s}$ ($s<i$) because those have weighted degrees either $d_{2}-d_{1}$ or $3d_{1}$, both different from 0. For Case 2, all modified weighted degrees in $I_{0}$ become negative; so $I_{0}$-vertices cannot re-enter $I_{0,0}$.
After handling $x_{i}$, update $I_{0,0},X_{2^{+}}$ and $X^{\ast}_{2^{+}}$, respectively. Note that $x_{i}$ and all previously processed vertices $x_{j}$ ($j<i$) are no longer in $X^{\ast}_{2^{+}}$ (no longer in $X_{2^{+}}$ either), since they either have weighted degree $d_{2}-d_{1}$ (which differs from 0), or have weighted degree $3d_{1}$ while all their neighbors in $I_{0}$ have weighted degrees different from $3d_{1}$. If $X^{\ast}_{2^{+}}=\emptyset$, we are done. If $X^{\ast}_{2^{+}}\neq\emptyset$ let $x_{i+1}\in X^{\ast}_{2^{+}}$ and the assumptions (i)-(iii) hold for $i+1$. Proceeding and update $X^{\ast}_{2^{+}}$ repeatedly, we eventually eliminate all conflicts in $G[I_{0},X_{2^{+}}]$. Moreover, conditions (1)-(3) remain valid currently. Hence no conflict remains incident to $X_{2^{+}}$ finally.
Remark on the dynamics of $X^{\ast}_{2^{+}}$: Since (in Case 1) $I_{0,0}$ may expand, $X^{\ast}_{2^{+}}$ may also expand. However, $X^{\ast}_{2^{+}}$ is included in $X_{2^{+}}$, and $X_{2^{+}}$ decreases at each step. So the recursive process will terminate with $X^{\ast}_{2^{+}}=\emptyset$.
2. Resolving conflicts incident to $X_{1}$.
After resolving conflicts for $X_{2^{+}}$, we update $I_{0,0}$. Note that $X_{1}$ remains unchanged. Let $X^{\ast}_{1}\subseteq X_{1}$ be the set of vertices in $X_{1}$ that have a neighbor in (updated) $I_{0,0}$. If $X^{\ast}_{1}=\emptyset$, we are finished.
Assume $X^{\ast}_{1}\neq\emptyset$ and let $M^{\ast}_{1}$ be the subset of $M_{1}$ consisting of edges incident to vertices in $X^{\ast}_{1}$. Assign weight $d_{1}$ to every edge in $M^{\ast}_{1}$. After the assignment, for each $x\in X^{\ast}_{1}$, $d_{w}(x)=-d_{1}$, while $d_{w}(y_{x})$ remains 0 and $d_{w}(z_{x})$ decreases by $d_{1}$, making $z_{x}$ a Type III vertex. This implies that condition (5) holds, and conditions (1)-(3) remain valid. Note that, at present, both $X^{\ast}_{2^{+}}$ and $X^{\ast}_{1}$ are empty, and so condition (6) holds. Condition (4) holds because: vertices with $d_{w}(x)=-3d_{1}$ come from Case 2. In Case 2.1, $x\in X_{3^{+}}$ has at least three neighbors in $I_{0}$; in Case 2.2, $x\in X_{2}$ has exactly two neighbors in $I_{0}$. In both cases, all neighbors $y$ of $x$ in $I_{0}$ satisfy $d_{w}(y)\neq-3d_{1}$ by construction. Thus, conditions (1)-(6) all hold, and $w$ is a proper $\{-d_{1},0,d_{2}\}$-edge weighting of $G$.
This completes the proof of Theorem 1. ∎
3. Concluding Remarks
This paper establishes that every regular graph is universally 3-edge-weightable. The conflict-resolution scheme introduced in this work may be applicable to other weighting and labeling problems. The two-phase approach–comprising global matching adjustments followed by local conflict resolution–provides a framework for handling parameter-dependent constructions in combinatorial graph theory.
Future work. Our result naturally leads to several open questions:
  • Does universal 3-edge-weightability extend to all nice graphs? (Problem 1.1)
  • What other graph classes (beyond regular graphs) admit this property?
  • Can the techniques be extended to study edge-weight choosability problems?
Acknowledgments
The first author is supported by NSFC (No. 11701195) and Fundamental Research Funds for the Central Universities (No. ZQN-904).